OpenStudy (anonymous):

Divide 3/10 by 2/5 Express your answer as a fraction in lowest terms.

OpenStudy (anonymous):

It'll look like \[\frac{ 3 }{ 10 } \times \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

Schen is right, when you are dividing two fractions, you take the second one, flip it and multiply.

OpenStudy (anonymous):

Now just multiply them to get \[\frac{ 15 }{ 20 }\]

OpenStudy (anonymous):

and then you can simplify that.

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

Yes, correct =)

OpenStudy (anonymous):

so the answer is 3/4?

OpenStudy (anonymous):

Yes, thats the lowest the fraction can be simplified into, so 3/4 will be the answer.

OpenStudy (anonymous):

thank you :)